# USTC-ICTS/PCFT-25-37

### v1

@@ Appendix A The topsector of sunset
$\left.+2\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (III-23) is
$-{f_{+++}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}-{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (36)
$-{2m_{1}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
$=$ $\left\{(7-2D){\partial\over\partial\eta_{3}}G_{111}-(D-3){\partial\over\partial\eta_{2}}G_{111}+2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.+(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}+2\eta_{1}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.

### v2

@@ Appendix A The topsector of sunset
$\left(\begin{array}[]{c| c |c |c| c| l}{\partial\over\partial\eta_{1}}&{\partial\over\partial\eta_{2}}&{\partial\over\partial\eta_{3}}&{\partial\over\partial\eta_{4}}&{\partial\over\partial\eta_{5}}&\\ \hline\cr&&&\widehat{O}_{0;I}&&{\rm(I-2)}+{\rm(I-3)}\\ \hline\cr&&&&\widehat{O}_{0;I}&{\rm(I-1)}-{\rm(I-3)}\\ \end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(54)
with
$\widehat{O}_{0;I}$ $=$ $=6-3D-2\eta_{4}{\partial\over\partial\eta_{4}}-2\eta_{5}{\partial\over\partial\eta_{5}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(55)
Solving them, we get $2$ T1B-type reduction rules for degree-1 operators ${\partial\over\partial\eta_{a}},a=4,5$. From the fourth and fifth rows, we can find $2$ T2B-type reduction rules, i.e., using $\partial\over\partial\eta_{3}$} to solve ${\partial\over\partial\eta_{i}},i=1,2$.
Checking module of second round for massless case: With new found reduction rules, the irreducible lattice points are $(0,0,n_{3},0,0)$, thus we need to go to the third round of computation.
Third round computation of massless case: Since only $n_{3}$ is not fully reduced, we can only act $\partial\over\partial\eta_{3}$} on $5$ reduction rules found in the second round of computation. The action on the reduction rule of $\partial\over\partial\eta_{4}$}, after simplification, gives
$0$ $=$ $\left\{\widehat{O}_{0;II}{\partial\over\partial\eta_{3}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$(56)
with
$\widehat{O}_{0;II}={(D-4)K^{2}\over 2s_{0}}-{K^{2}\over s_{0}}\eta_{3}{\partial\over\partial\eta_{3}}\penalty 10000\ \penalty 10000\ \penalty 10000\$(57)
where ... is the part of lower degree and boundary. Using the result (56), one can see that all nonzero $n_{3}$ can be reduced to $n_{3}=0$. So finally the only point can not be reduced using the reduction rule is the point $(0,0,0,0,0)$, which is the master integral for this case. Now we find the complete reduction rules.
Solving module of second round for massive case: There are $6$ degree-2 equations with only $5$ of them to be independent. The coefficient matrix is
$\left(\begin{array}[]{c |c |c |c|c|c|l}[1,1]&[2,2]&[3,3]&[1,2]&[1,3]&[2,3]&\\ \hline\cr&{-2m_{2}^{2}\over s_{0}}&&{-f_{+++}\over s_{0}}&{-2m_{3}^{2}\over s_{0}}&{-2m_{3}^{2}\over s_{0}}&{\rm(II-12)}\\ \hline\cr&&{-2m_{3}^{2}\over s_{0}}&{-2m_{2}^{2}\over s_{0}}&{-f_{+++}\over s_{0}}&{-2m_{2}^{2}\over s_{0}}&{\rm(II-13)}\\ \hline\cr{2m_{1}^{2}\over s_{0}}&&&{f_{+++}\over s_{0}}&{2m_{3}^{2}\over s_{0}}&{2m_{3}^{2}\over s_{0}}&{\rm(III-12)}\\ \hline\cr&&{2m_{3}^{2}\over s_{0}}&{2m_{1}^{2}\over s_{0}}&{2m_{1}^{2}\over s_{0}}&{f_{+++}\over s_{0}}&{\rm(III-23)}\\ \hline\cr&{-2m_{2}^{2}\over s_{0}}&{2m_{3}^{2}\over s_{0}}&&&&{\rm(II-23)}\\ \hline\cr{2m_{1}^{2}\over s_{0}}&&{-2m_{3}^{2}\over s_{0}}&&&&{\rm(III-13)}\end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(65)
The coefficient matrix of $3$ degree-1 equations is